OpenStudy (anonymous):

A silversmith mixed 45 g of an 80% silver alloy with 50 g of a 25% silver alloy. What is the percent concentration of the resulting alloy? (Round your answer to one decimal place.)

OpenStudy (anonymous):

@zepdrix if you're available can you help me out ty if so!

zepdrix (zepdrix):

Hmmmm let's see if i can explain this in a way that makes sense D: These concentration problems always get me a bit tripped up.

zepdrix (zepdrix):

`45 grams at 80% concentration` + `50 grams at 25% concentration` = `the 45 + 50 at some new concentration`. \[\Large\bf\sf 45\cdot 80\text{%}+50\cdot 25\text{%}\quad=\quad 95\cdot X\text{%}\]

zepdrix (zepdrix):

We'll write our percents as decimals and try to solve for x.\[\Large\bf\sf 45\cdot0.8+50\cdot 0.25\quad=\quad 95\cdot \frac{x}{100}\]

zepdrix (zepdrix):

Mmmm am I doing this right maybe? :d Is the first step kinda confusing?

OpenStudy (anonymous):

i have everything like that except the last part (x/100) i have =95x

OpenStudy (anonymous):

I have 0.8(45)+.25(50)=95x

zepdrix (zepdrix):

Ummm ok that's fine. Just realize that your `x` that you end up with is not a percent. You'll need to convert it to a percent afterwards.

OpenStudy (anonymous):

I though the answer was either 51 or 52% rounding but that is incorrect

zepdrix (zepdrix):

51% was incorrect? D: Hmm

zepdrix (zepdrix):

Oh one decimal, so we should get ummmm 51.1%

zepdrix (zepdrix):

Maybe that will make a difference? :o

OpenStudy (anonymous):

Thumbs up! TY!! @zepdrix

zepdrix (zepdrix):

Oh was it that simple? :) yay team!